OpenStudy (anonymous):

Given the accerleration, initial position of a body moving along a coordinate line at time t, find the bodys position at time t a=120, v(0)=2,s(0)=-4 A s=60t^2 +2t-4 B s=120t^2 +2t-4 Cs= -60t^2-2t-4 D s=60t^ =2t Can someone help

OpenStudy (lgbasallote):

is this the whole question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its Cal 1

OpenStudy (lgbasallote):

i believe you can use the formula \[s_2 - s_1 = v_1 t + \frac 12 at^2\] where s2 is final position s1 is initial position v1 is inital speed t = time a = acceleration

OpenStudy (lgbasallote):

plug in your values and tell me what you ahve

OpenStudy (lgbasallote):

have*

OpenStudy (anonymous):

i have 60t^2 + 2t is that right. But what do i do with the s(0)=-4?

OpenStudy (lgbasallote):

hmm..if you tried plugging in the values into the formula i stated you would have gotten s - (-4) = 2t + 1/2 (120) t^2 do you see it now?

OpenStudy (anonymous):

so would it be s=60t^2 +2t -4?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Thanks so much i appreciate your help!

OpenStudy (lgbasallote):

you're welcome :DD

OpenStudy (anonymous):

hey do you think you could help me with another ?

OpenStudy (lgbasallote):

why not :D

OpenStudy (anonymous):

Find the value or values of c that satisfy the equation f(b)-f(a)/b-a= f'(c) in the conclusion of the Mean Value Theroem for the function and interval. Round to the nearest thousandth F(x)=tan^-1(x), [-sqrt(3), sqrt(3)]

OpenStudy (lgbasallote):

awww not good with this sorry :(

OpenStudy (lgbasallote):

i hated those theorems lol

OpenStudy (anonymous):

its cool thank anyway i hate them too! lol